OpenStudy (anonymous):

okay so i have to find x on a complementary angle it is 6x and 5x-20 can you explain please?

OpenStudy (asnaseer):

if two angles are complementary, then they add up to 90 degrees so if one of your angles is 6x and the other is 5x-20, then we have: 6x + 5x - 20 = 90

OpenStudy (asnaseer):

you just need to solve this to find x

OpenStudy (asnaseer):

do you understand?

OpenStudy (anonymous):

trying to so i would actually take the 20 from the 90? and then try to find the answer?

OpenStudy (asnaseer):

not quite. you need to ADD 20 to both sides of the equation.

OpenStudy (asnaseer):

so you would get: 6x + 5x - 20 + 20 = 90 + 20 6x + 5x = 110

OpenStudy (anonymous):

so x is 10 right?

OpenStudy (asnaseer):

perfect - well done!

OpenStudy (anonymous):

ty you made it easy to understand

OpenStudy (asnaseer):

I'm glad I could help :-)